Show that the tangents from the point with position vector −2i − 3j to the ellipse
4x
2 + 9y
2 = 36 are perpendicular.
1
Expert's answer
2020-03-13T15:10:42-0400

Equation of ellipse is given as


"4x^2 + 9y^2 = 36" , the general form of equation is


"\frac{x^2}{9} +\frac{y^2}{4} = 1"


on comparing with general equation of ellipse


"\frac{x^2}{a^2}+\frac{y^2}{b^2}= 1"


here, a= 3 , and b=2,

here the position vector of point is given as -2i-3j so the coordinate of point will be (-2,-3)


Now, we know that equation of tangent to the given ellipse with m as slope will be


"y= mx+\sqrt{a^2m^2+ b^2}"


"y= mx+ \sqrt{9m^2+4}"


and it passes through the point (-2,-3)


so the equation become

"-3=-2m+\sqrt{9m^2+4}"


"2m-3=\sqrt{9m^2+4}"


now squaring the both sides we get


"(2m-3)^2= (\sqrt{9m^2+4})^2"


"4m^2+9-12m= 9m^2+4"


"5m^2+12m-5=0"


this is quadratic equation we can find the value of m by solving this equation using quadratic formula,

"m_1=\frac{-12 +\sqrt{144+100}}{10} ,m_2=\frac{-12 -\sqrt{144+100}}{10}" .


These are the slopes of two tangents from the ellipse passing through the point(-2,-3),


now both will be perpendicular is


"m_1m_2=-1"


LHS"=" "\frac{-12 +\sqrt{144+100}}{10} \times \frac{-12 -\sqrt{144+100}}{10}"

LHS=-1 = RHS


so the tangents are perpendicular


Hence proved.........